I am trying so very hard to get this correct, as well as place it in an equation that makes sense. Each time I try, I fumble. What I need is to see it before my eyes, completely. I am so visually oriented and I feel the only way I'm going to get it is by looking at it. Thanks. Here's the problem:
3. Jason had to submit five animation projects as per his syllabus in the final semesters. He scored 82, 85, 78, and 88 points out of 100 in four projects. What should be his score in the fifth project so that the average of all the projects is at least 85?
: I am trying so very hard to get this correct, as well as place it in an equation that makes sense. Each time I try, I fumble. What I need is to see it before my eyes, completely. I am so visually oriented and I feel the only way I'm going to get it is by looking at it. Thanks. Here's the problem:
3. Jason had to submit five animation projects as per his syllabus in the final semesters. He scored 82, 85, 78, and 88 points out of 100 in four projects. What should be his score in the fifth project so that the average of all the projects is at least 85?

Answer by checkley77(3848)

(82+85+78+88+x)/5>=85
(33+x)/5>=85
33+x>=5*85
33+x>=425
x>=425-333
x>=92 is the minimum grade on the last test to maintain at least an 85 average.